# Nonlinear anisotropic equilibrium reconstruction in axisymmetric magnetic mirrors

### v3

@@ IV Conclusion and future work
In addition to theoretical and experimental improvements above, several computational improvements will be considered in the future. The MORBO multiobjective optimization using the algorithm in Daulton 2022Daulton et al. (2022) will be implemented in which each diagnostic signal is treated as an individual objective function by the reconstruction algorithm. This approach will generate a pareto front of solutions that will allow even more effective quantification of uncertainty. Reconstructions will also be modified include the impact of parallel flows that could potentially be relevant outside the core in the mirror expanders where there are low fields and fast outflows. These flows are unlikely to affect core reconstructions but could affect the field line curvature and may be important for calculations of stability. Finally, while WHAM has a large vacuum boundary that allows us to neglect wall currents, future magnetic mirrors are anticipated to have more conformal walls as they can be beneficial for stability. Therefore, the impact of the wall and its 3D structure on flux measurements will need to be considered. More sophisticated equilibrium solvers, like the Tokamaker codeHansen et al. (2024) and 3D MHD solvers, which can account for resistive walls more accurately should be adapted for mirrors.
@@ Appendix A Derivation of the model for $f_{\mathrm{hot}}(v,\theta)$
is the slowing down time Cordey (1976), $\ln\Lambda$ is the Coloumb logarithm, $\epsilon_{0}$ is the permittivity of free space, $m_{i}$ and $m_{e}$ are the ion and electron masses, $Z_{i}$ is the ion charge, $e$ is the fundamental charge, $T_{e}$ is the electron temperature, and $n_{e}$ is the electron density.

### v4

@@ IV Conclusion and future work
In addition to theoretical and experimental improvements above, several computational improvements will be considered in the future. The MORBO multiobjective optimization using the algorithm in Daulton 2022Daulton et al. (2022) will be implemented in which each diagnostic signal is treated as an individual objective function by the reconstruction algorithm. This approach will generate a pareto front of solutions that will allow even more effective quantification of uncertainty. Reconstructions will also be modified to include the impact of flows, as in some cases with relatively cold plasmas, rotational velocities can become large enough to violate $M^{2}\ll 1$. It is also possible that parallel flows could be relevant outside the core in the mirror expanders where there are low fields and fast outflows. These flows are unlikely to affect core reconstructions but could affect the field line curvature and may be important for calculations of MHD stability. Finally, while WHAM has a large vacuum boundary that allows us to neglect wall currents, future magnetic mirrors are anticipated to have more conformal walls as they can be beneficial for stability. Therefore, the impact of the wall and its 3D structure on flux measurements will need to be considered. More sophisticated equilibrium solvers, like the Tokamaker codeHansen et al. (2024) and 3D MHD solvers, which can account for resistive walls more accurately should be adapted for mirrors.
@@ Appendix A Derivation of the model for $f_{\mathrm{hot}}(v,\theta)$
is the slowing down time Cordey (1976), $\ln\Lambda$ is the Coulomb logarithm, $\epsilon_{0}$ is the permittivity of free space, $m_{i}$ and $m_{e}$ are the ion and electron masses, $Z_{i}$ is the ion charge, $e$ is the fundamental charge, $T_{e}$ is the electron temperature, and $n_{e}$ is the electron density.